OpenStudy (anonymous):

Given the function f defined by f(x)=x^2-2x-3, find f'(3)

OpenStudy (anonymous):

We would need our derivative. So f'(x) = 2x - 2. Now plugging in 3 we get f'(3) = 2(3) - 2 or 4. So f'(3) = 4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Thanks, I understand it now. Just take the derivative and then plug in 3.

OpenStudy (bahrom7893):

YAY

OpenStudy (anonymous):

f'(x)=2x-2 f'(3)=2(3)-2=4